Please help me solve 3h^2 + 11h + 6
Answer by stanbon(75887)

Please help me solve 3h^2 + 11h + 6 = 0
---
AC Method
A*C = 3*6 = 18 = 2*9
B = 11 = 2 + 9
------
Rewrite the equation::
3h^2 + 2h+9h + 6 = 0
----
Factor in groups::
h(3h+2) + 3(h+2) = 0
Factor::
(3h+2)(h+3) = 0
------
Ans: h = -2/3 or x = -3
-----------------
Cheers,
Stan H.
----------